\begin{thebibliography}{10}

\bibitem{LongZheng}
Longyun Ding and Yang Zheng.
\newblock On equivalence relations induced by {P}olish groups.
\newblock {\em 2204.04594}, 2022.

\bibitem{OPENPROBLEMS}
Jerome~Buzzi et~al.
\newblock Open problems in descriptive set theory and dynamical systems.
\newblock {\em Arxiv 2305.00248}, 2023.

\bibitem{Foremanreduction}
Matthew Foreman.
\newblock What is a {B}orel reduction?
\newblock {\em Notices Amer. Math. Soc. 65, no. 10, 1263–1268.}, 2018.

\bibitem{Surveyforeman}
Matthew Foreman.
\newblock The {C}omplexity and the {S}tructure and {C}lassification of {D}ynamical {S}ystems.
\newblock {\em Encyclopedia of Complexity and Systems Science, Springer}, 2022.

\bibitem{Foremanweiss}
Matthew Foreman and Benjamin Weiss.
\newblock An anti-classification theorem for ergodic measure preserving transformations.
\newblock {\em Journal of the European Mathematical Society 6 (3), 277-292}, 2004.

\bibitem{SmoothMPTnonBorel}
Matthew Foreman and Benjamin Weiss.
\newblock Measure preserving diffeomorphisms of the torus are unclassifiable.
\newblock {\em J. Eur. Math. Soc. 24 (8): 2605–2690}, 2022.

\bibitem{Kakutaninonborel}
Marlies Gerber and Phillip Kunde.
\newblock Non-classifiability of ergodic flows up to time change.
\newblock {\em Inventiones Mathematicae, 239, pp.527-619}, 2025.

\bibitem{BorelEQ}
Vladimir Kanovei.
\newblock {\em Borel Equivalence Relations : Structure and Classification}.
\newblock University Lecture Series. American Mathematical Society, 2008.

\bibitem{WeakliymixingnonBorel}
Phillip Kunde.
\newblock Anti-classification results for weakly mixing diffeomorphisms.
\newblock {\em Mathematische Annalen, Volume 390, pp. 5607–5668}, 2024.

\bibitem{Foremanlouveaubook}
Alain~Louveau Matthew~Foreman, Alexander S.~Kechris and Benjamin Weiss.
\newblock {\em Descriptive Set Theory and Dynamical systems}.
\newblock Cambridge University Press, 2000.

\bibitem{NonBorelnessEMPT}
Daniel J.~Rudolph Matthew~Foreman and Benjamin Weiss.
\newblock The conjugacy problem in ergoic theory.
\newblock {\em Annals of Mathematics 173, pp. 1529-1586}, 2011.

\bibitem{entropy}
D.S. Ornstein.
\newblock Some new results in the kolmogorov–sinai theory of entropy and ergodic theory.
\newblock {\em Bull. Amer. Math. Soc. 77, 878–890}, 1971.

\bibitem{Unblanced}
Aristotelis Panagiotopoulos and Shaun Allison.
\newblock Dynamical obstructions to classification by (co)homology and other tsi-group invariants.
\newblock {\em Transactions of the American Mathematical Society}, 2021.

\bibitem{Choquetcomplete}
Marcin Sabok.
\newblock Completeness of the isomorphism problem for separable c*-algebras.
\newblock {\em Inventiones Mathematicae, Volume 204, pages 833–868}, 2016.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb}

\usepackage{comment}
\usepackage{color}
\usepackage{mathrsfs}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{fact}[theorem]{Fact}
\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{ques}[theorem]{Question}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{conjecture}[theorem]{Conjecture}
\theoremstyle{remark}

\numberwithin{equation}{section}

\begin{document}

\title[Unbalancedness of ergodic measure-preserving transformations]{Unbalancedness of the conjugacy relation of ergodic measure-preserving transformations}


\author{Bo Peng}
\address{Department of Mathmatics and Statistics, McGill University. 805 Sherbrooke Street West Montreal, Quebec, Canada, H3A 2K6}
\email{bo.peng3@mail.mcgill.ca}


\date{}

\dedicatory{}

\commby{}


\maketitle
\begin{abstract}
    We show that the isomorphism of ergodic measure-preserving transformations is not Borel reducible to the relation induced by the conjugacy action of the full group of an ergodic measure-preserving transformation on itself. This answers a question of Le Ma\^{i}tre in the negative and gives a positive indication towards a conjecture of Sabok. In fact, we prove that the isomorphism of ergodic measure-preserving transformation is unbalanced, which answers another question of Le Ma\^{i}tre in the positive.
\end{abstract}
\section{introduction}

The history of classifying measure-preserving transformations (MPT) up to conjugacy can be traced back to the early 20th century when von Neumann proved that ergodic transformations with discrete spectrum can be classified by the eigenvalues of the associated Koopman operator. In 1971, Ornstein \cite{entropy} proved that Bernoulli shifts are classified by their entropy. Despite many efforts and partial results, the general problem of classifying MPTs was open for quite a long time. 

Let $E$ and $F$ be two equivalence relations on Polish spaces $X$ and $Y$, respectively. A Borel function from $X$ to $Y$ is called a \textbf{Borel reduction} from $E$ to $F$, if for any $x_1,x_2\in X$, we have
$$
x_1Ex_2\,\,\mbox{if and only if}\,\,f(x_1)Ff(x_2).
$$
We say $E$ is \textbf{Borel reducible} to $F$ if there is a Borel reduction from $E$ to $F$ and denote it by $E\leq_B F$. For more details and motivation of a Borel reduction see \cite{Foremanreduction}.  By applying this theory to the classification problems, one can prove impossibility of various types of classification. We call such a result an \textbf{anti-classification theorem}.

In the last 20 years, many types of anti-classification results for isomorphism of MPTs were proved. Foreman and Weiss \cite{Foremanweiss} proved that the conjugacy action of MPTs on ergodic measure-preserving transformations (EMPT) is \textbf{turbulent}. This shows that von Neumann's classification of EMPTs with discrete spectrum is impossible to generalize to classify all EMPTs. A great accomplishment in this direction is that Foreman, Rudolph and Weiss \cite{NonBorelnessEMPT} proved that the isomorphism relation of EMPTs is not Borel. By developing completely new techniques, Foreman and Weiss \cite{SmoothMPTnonBorel} also proved this for smooth ergodic measure-preserving diffeomorphisms of the on the torus, disk and annulus. Recently, Gerber and Kunde \cite{Kakutaninonborel} showed that the Kakutani equivalence for MPTs is not Borel and Kunde \cite{WeakliymixingnonBorel} proved non-Borelness of weakly mixing diffeomorphisms on the torus, disk and annulus, for both conjugacy and Kakutani equivalence. 

However, the exact complexity of the isomorphism of ergodic measure-preserving transformations remains open. A famous result of Sabok \cite{Choquetcomplete} shows that the affine homeomorphism relation of metrizable compact Choquet simplices is a complete orbit equivalence relation. Since the space of invariant measures of a homeomorphism can be equipped with a Choquet simplex structure, Sabok made the following conjecture (See \cite[Open problem 12]{Surveyforeman} and \cite[Conjecture 4.1]{OPENPROBLEMS}):

\begin{conjecture}
    (Sabok Conjecture) The isomorphism relation of EMPTs is a complete orbit equivalence relation.
\end{conjecture}

However, in the same survey \cite{OPENPROBLEMS}, Le Ma\^{i}tre asked the following question. Let $T\in {\rm Aut}(X,\mu)$ be a measure-preserving transformation. We define its \textbf{full group} $[T]$ as
$$
[T]=\{U\in {\rm Aut}(X,\mu)| \forall x\,\, U(x)\in \{T^n(x)\}_{n\in \mathbb{Z}}\}.
$$
Fix an EMPT $T_0$.
\begin{ques} \label{Q0}
    (Le Ma\^{i}tre \cite[Question 4.2]{OPENPROBLEMS}) Does the conjugacy on ${\rm Aut}(X,\mu)$ Borel reduce to the conjugacy relation on $[T_0]$? Does it when we restrict to ergodic transformations?
\end{ques}

Note that Sabok's conjecture implies the negative answer to the above question. The reason is that the Halmos metric is a complete bi-invariant metric on $[T_0]$. A Polish group with a complete two-sided invariant metric is called a \textbf{TSI Polish group}. It is known (see \cite{Unblanced}) that any action induced by this kind of group can not be a complete orbit equivalence relation.

In \cite{Unblanced}, Allison and Panagiotopoulos developed a new theory of \textbf{generic unbalancedness} to prove that an equivalence can not be Borel reducible to any ${\rm TSI}$ group action. See also \cite{LongZheng} for some applications and examples of this theory.

\begin{definition}
    Let $G$ be a Polish group acting in a Borel way on a Polish space $X$. Let $x,y\in X$, we write $x\leftrightsquigarrow y$ if for every open neighborhood $V$ of the identity of $G$ and every open set $U\subset X$ having non-empty intersection with the orbit of $x$ or $y$, there exist $g^x,g^y\in G$ with $g^xx\in U$, $g^yy\in U$, so that:
    $$
         g^yy\in \overline{V(g^xx)}\,\,\mbox{and}\,\,(g^xx)\in \overline{V(g^yy)}.
    $$
    
\end{definition}

    Note that if $x \leftrightsquigarrow y$, then $\overline{Gx}=\overline{Gy}$ (see \cite{Unblanced}). Also, for any $y\in X$ sharing the same orbit with $x$, we have $x\leftrightsquigarrow y$.

\begin{definition}
   \cite[Definition 1.1]{Unblanced} Let $X$ be a Polish $G$-space and let $C\subset X$ be a $G$-invariant set. The \textbf{unbalanced graph} associated to the action of $G$ is the graph $(C/G, \leftrightsquigarrow)$ where $C/G=\{[x]|x\in C\}$ and $[x]\leftrightsquigarrow [y]$ if and only if $x \leftrightsquigarrow y$.
\end{definition}

\begin{definition}
    \cite{Unblanced} The group action $G$ on $X$ is \textbf{unbalanced} if it has meager orbits and for every $G$-invariant comeager $C\subset X$, there is a comeager $D\subset C$ so that for all $x,y\in D$ there is a path between $[x]$ and $[y]$ in $C/G$.
\end{definition}

 In other words, the action of $G$ on $X$ is unbalanced if for every $G$-invariant subset $C$, there is a comeager $D\subset C$ such that for any $x,y\in D$ there is a sequence $x_0,x_1,\dots,x_n \in C$ with $x_0=x$ and $x_n=y$ such that $x_{i-1} \leftrightsquigarrow x_i$ for all $0<i<n$. 

\begin{theorem} \label{unbalance}
    {\rm (Allison and Panagiotopoulos \cite{Unblanced})} If the Polish $G$-space is generically unbalanced, then the orbit equivalence relation is not classifiable by ${\rm TSI}$ Polish group actions. 
\end{theorem}

In \cite[Question 4.3]{OPENPROBLEMS}, Le Ma\^{i}tre also asked the following question:

\begin{ques} \label{Q1}
    (Le Ma\^{i}tre \cite[Question 4.3]{OPENPROBLEMS}) Is the conjugacy action of ${\rm MPTs}$ on itself unbalanced? Is it when we restricted to ${\rm EMPTs}$? 
\end{ques}

By the result of \cite{Unblanced}, a positive answer to Question \ref{Q1} implies the negative answer to Question \ref{Q0}. In this paper, we answer Question \ref{Q1} in the positive and thus Question \ref{Q0} in the negative.

\begin{theorem} \label{Mainthm}
    The conjugacy actions of ${\rm Aut}(X,\mu)$ on both itself and on ergodic measure-preserving transformations are unbalanced.
\end{theorem}

\begin{corollary}\label{cor}
    The conjugacy relation of any generic subset of ${\rm Aut}(X,\mu)$ and on ergodic transformations is not classifiable by any ${\rm TSI}$ group actions. 
\end{corollary}

In fact, in Section \ref{shortcut}, we first give a direct argument to answer Question \ref{Q0} and actually this shows the anti-classification theorem even holds when restricted to ${\rm EMPTs}$ with discrete spectrum. In Section \ref{Mainresult}, we give a proof of Theorem \ref{Mainthm}.

\section{A direct proof of the anti-classification} \label{shortcut}
  Recall that $=^+$ is the equivalence relation defined on all sequences of reals as follows,  for all $(x_n),(y_n)$,
$$
(x_n)=^+(y_n)\,\,\mbox{if}\,\, \exists\,g\in S_{\infty}\,\forall n\,x_n=y_{g(n)}.
$$

We will use the following theorem of Foreman and Louveau \cite{Foremanlouveaubook}. 

\begin{theorem} \label{FLTHM}
    {\rm (Foreman and Louveau \cite[Theorem 65]{Foremanlouveaubook})} $=^+$ is Borel bireducible with the isomorphism  relation of the ergodic transformation with discrete spectrum.
\end{theorem}

Below we give a direct proof of Corollary \ref{cor}. In fact, we give two proofs. The first one uses the machinery of unbalancedness of Allison and Panagiotopoulos, it essentially follows the work of \cite{Unblanced}, but we prove it directly. The second proof uses pinned equivalence relation.


\begin{proof}[Proof of Corollary \ref{cor}]

    By Theorem \ref{FLTHM} it is enough to show $=^+$ is not Borel reducible to any ${\rm TSI}$ Polish group action since the group $[T_0]$ is a TSI group.
    
    Let $D$ be the set of all countable dense sequences of  reals. Note that $D$ is a dense $G_{\delta}$ subset of $\mathbb{R}^{\omega}$.

    We will prove the following:

    \begin{lemma}\label{unb}
        The action of $S_{\infty}$ on $D$ is unbalanced.
    \end{lemma}
  \begin{proof}
  
   By the choice of $D$, every orbit is dense. Next we prove that every orbit is meager. Let $x\in D$ and let 
    $$
    B_x=\{b\in D\mid \,\, \forall m\,b(1)\neq x(m) \}.
    $$
     Note that $B_x=((\mathbb{R}-\{x(m)\}_m) \times\mathbb{R}^{\omega}) \cap D$ which is the intersection of a product of two comeagers and  $D$. Thus, $B_x$ is comeager. Since we remove all $\{x_m\}$ in the first coordinate, we have that $B_x$ is not in the orbit of $x$. Thus, $B_x$ is a subset of the complement of the orbit of $x$ which implies that the orbit of $x$ is meager.

     Now we prove that for any $x,y\in D$ there is a path on the unbalanced graph connecting $x$ and $y$. 

    \begin{lemma}\label{keylemma}
        For any $x,y\in D$, if $x,y$ share a dense infinite subset, in other words, if there exists two sequences of natural numbers $\{p_n\},\{q_n\}$ such that $\{x_{p_n}\}=\{y_{q_n}\}$ is dense in $\mathbb{R}$, then $x \leftrightsquigarrow y$.
    \end{lemma}
    \begin{proof}
        Assume that both $x$ and $y$ contain the dense sequence $\{d_n\}$.
         
        Since every orbit is dense, every open set intersects both orbits of $x$ and $y$, fix an open set $U \subset C$. Take an any open neighborhood of $1_{S_{\infty}}$, say $V$. We may assume that $V$ is the set of those permutations which fix the first $k$ many elements. Since $\{d_m\}$ is dense, we can find $d_{t_1},\dots,d_{t_r}$ such that $r\geq k$ and all elements with the first $r$ coordinates equal to $(d_{t_1},\dots,d_{t_r})$ are all in $U$. Then take $g^x$ and $g^y$ to be two elements in $S_{\infty}$ such that the first $r$-elements of $g^xx$ and $g^yy$ are the same and equal to $(d_{t_1},\dots,d_{t_r})$. So we have $g^xx,g^yy\in U$. Now since $V$ fixes the first $k$ many elements, $k\leq r$ and $x$ enumerates a dense sequence, we have that $\overline{Vg^xx}$ is the set of all sequences whose first $k$-coordinate is equal to $(d_{t_1},d_1,\dots,d_{t_k})$.  This implies that $\overline{Vg^xx}=\overline{Vg^yy}$ which means $x \leftrightsquigarrow y$.
    \end{proof}

    For all $x,y\in D$, let $z$ be a sequence such that $z(2n)=x(n)$ and $z(2n+1)=y(n)$. We have $z\in D$, and by Lemma \ref{keylemma}, we know that $x\leftrightsquigarrow z$ and $z\leftrightsquigarrow y$.

\end{proof}
Since $D$ is $S_{\infty}$-invariant, we know that $E^D_{S_{\infty}}$ is Borel reducible to $=^+$. 

An alternative proof goes as follows.  $=^+$ is unpinned \cite[Claim 17.2.1]{BorelEQ}, ${\rm CLI}$ group actions are all pinned \cite[Chapter 17.4]{BorelEQ} and ${\rm TSI}$-groups are all ${\rm CLI}$. Thus, $=^+$ is not reducible to any ${\rm TSI}$ group actions.

\end{proof}

\section{Global unbalancedness} \label{Mainresult}

We use the same notation as in \cite{Foremanweiss}.  ${\rm Aut}(X,\mu)$ can be equipped with a natural topology whose neighborhood basis is the following collection of sets: 
$$
\mathcal{N}_T(A_1,A_2,..,A_n, \epsilon)=\{S\in {\rm Aut}(X,\mu)| \forall i\leq n\, \mu(S(A_i)\Delta T(A_i))<\epsilon)\}
$$
where $T\in {\rm Aut}(X,\mu)$ and $A_i\subset X$ are pairwise disjoint measurable sets. 
Halmos defined a metric which is finer than the above topology: 
$$
d(S,T)=\mu(\{x\in X|S(x)\neq T(x)\}).
$$

For $T\in {\rm Aut}(X,\mu)$ and $M\in \mathbb{N}$, we call $T$ an \textbf{$M$-periodic} transformation if for a.e. $x\in X$, we have $T^Mx=x$ and for any $1\leq n<M$ we have $T^nx\neq x$. 
\begin{theorem}
    The conjugacy action of ${\rm Aut}(X,\mu)$ on the set of ergodic measure- preserving transformations is unbalanced.
\end{theorem}

\begin{proof}
    In \cite[Theorem 12]{Foremanweiss}, Foreman and Weiss proved that every orbit of this action is dense and meager. Thus, we only need to verify the last requirement of unbalancedness. Actually we are going to prove that for any two ergodic transformations $T_1$ and $T_2$, we have $T_1 \leftrightsquigarrow T_2$. Since ${\rm EMPTs}$ are generic in ${\rm MPTs}$, this implies that the whole conjugacy action of ${\rm MPTs}$ on itself is unbalanced.
\begin{lemma} \label{Rokhlin}
    {\rm (Rokhlin's Lemma)} Let $T$ be an ergodic measure-preserving transformation on $X$, $\epsilon>0$. Then there is a set $B\subset X$ such that
    \begin{enumerate}
        \item $B,T\,B,T^2B,\dots,T^{M-1}B$ are pairwise disjoint.
        \item The measure of $\bigcup_{i<M}T^iB$ is at least $1-\epsilon$.
    \end{enumerate}
\end{lemma}

We will call $B,T\,B,T^2B,\dots,T^{M-1}B$ a \textbf{$(M,\epsilon)$ Rokhlin tower with base $B$}. The following corollary is a well-known consequence of the Rokhlin's Lemma.

\begin{corollary} \label{Corouse}
    For any ergodic transformation $R\in {\rm Aut}(X,\mu)$ and $\eta>0$, there exists an $M_0$, such that for all $M\geq M_0$, there exists an $M$-periodic transformation $P$ such that $d(R,P)<\eta$.
\end{corollary}

\begin{proof}
   Fix $R$. Take $M_0$ such that $\frac{1}{M_0}<\frac{\eta}{2}$. Now for any $M\geq M_0$, by Lemma \ref{Rokhlin}, we can find a $(M, \frac{\eta}{2})$ Rokhlin tower with base $B$ for $R$. Then we can take any $M$-periodic MPT, say $P$, such that $P=R$ on $\bigcup_{i<M-1}R^i B$. Now since $R=P$ on $\bigcup_{i<M-1}R^i B$ and by the definition of Rokhlin tower, we know 
   $$
   \{x| R(x)\neq P(x)\} \subset R^{M-1} (B)\,\cup\,(X-\bigcup_{i<M}R^iB).
   $$
   Since 
   $$
   \mu (R^{M-1}B)+ \mu (X-\bigcup_{i<M}R^iB)< \frac{1}{M}+\frac{\eta}{2}<\eta,
   $$
   we have $d(R,P)<\eta$.
\end{proof}

To prove the unbalancedness, we need the following lemma from \cite{Foremanweiss}.
\begin{lemma} \label{keylemmaforeman}
    {\rm \cite[Lemma 9]{Foremanweiss}} Let $S,T$ be ergodic measure-preserving transformations and $\mathcal{W}$ a neighborhood of $S$. Then there is a $T'\in \mathcal{W}$ that has the same orbits almost everywhere as $T$.
\end{lemma}

The idea of the following lemma comes from \cite[Claim 14]{Foremanweiss},
\begin{lemma} \label{closeorbit}
     Fix $\epsilon>0$. Let $S$ and $T$ be two ergodic measure-preserving transformations, suppose $S$ has the same orbits as $T$ almost everywhere. If $d(S,T)<\epsilon$, then for any $\delta>0$, we can find $H$ such that $d(H,{\rm Id})<\epsilon$ and $d(HTH^{-1},S)<\delta$. 
\end{lemma}
\begin{proof}
    Since $S$ and $T$ are orbit equivalent, we can find an $M$ large enough such that the set
    $$
    L_0=\{x\in X| \{S(x),S^{-1}(x)\}\not \subset \{T^{-M+1}x, T^{-M+2}x,\dots, T^{M-1}x\}\}
    $$
    has measure less than $\frac{\delta}{3}$. 

    Now take a large enough $N$ such that $\frac{2M}{N}<\frac{\delta}{3}$, and choose an $(N,\frac{\delta}{3})$ Rokhlin tower with base $B$ for $T$. We let $L_1$ be the union of the first and the last $M$-levels of the tower. Then by the definition of Rokhlin tower, we know that $\mu(L_1)<\frac{\delta}{3}$. Let $L_2$ be the collection of points which are not in the Rokhlin tower. Thus we know the measure of the union $L=L_0\cup L_1 \cup L_2$ is less than $\delta$.

    Let $y\in B$. Then for all $x\in \{T^iy|M\leq i \leq N-M\}\setminus L_0$, both $Sx$ and $S^{-1}x$ belong to $\{T^iy:0\leq i\leq N-1\}$.

    Since both $T$ and $S$ are not periodic, we can choose a cyclic permutation $\sigma_y$ of $\{0,1,\dots, N-1\}$ such that for all $i,j<N-1$, if $S(T^i)=T^jy$ then $\sigma_y(i)=j$. For each permutation $\sigma$ of $\{0,\dots,N-1\}$, the collection $B_{\sigma}=\{y:\sigma_y=\sigma\}$ is measurable and the sets $B_{\sigma}$ partitions $B$.

    Given $\sigma$, a cyclic permutation of $\{0,1,\dots, N-1\}$, let $\tau_{\sigma}$ be the permutation of $\{0,1,\dots, N-1\}$ such that $\tau_{\sigma}\sigma \tau_{\sigma}^{-1}$ is the cyclic permutation $(0,1,\dots,N-1)$. Define $u_{\sigma}$ on $\bigcup_{i<N}T^iB_{\sigma}$ by setting
    $$
     u_{\sigma}(T^iy)=T^{\tau_{\sigma}(i)}y
    $$
    for $y\in B_{\sigma}$ and $u_{\sigma}(x)=x$ for $x\not \in \bigcup_{i<N}T^iB_{\sigma}$.

    Now let $H\in {\rm Aut}(X,\mu)$ be such that $H=u_{\sigma}$ on $\bigcup_{i<N}T^iB_{\sigma}$ and $H(x)=x$ otherwise.
    Since the collections of points moved by the various $u_{\sigma}$'s are pairwise disjoint, the $u_{\sigma}$'s commute and $H$ is well defined. 

     We claim that each $u_{\sigma}$ is identity on $\{x|T(x)=S(x)\}$. We choose $y\in B\setminus L$, suppose $T^i(y)\in \{x|T(x)=S(x)\}$, then $S(T^iy)=T^{i+1}y$.  Thus, we know that $\sigma(i)=i+1$. This implies that $\tau_{\sigma}(i)=i$. By the definition of $u_{\sigma}$, we know  $u_{\sigma}(T^iy)=u_{\sigma}(T^{\tau_{\sigma}(i)}(y))=T^iy$.
    
    This means, $H$ is identity on the set 
    $$
   \{x|T(x)=S(x)\}.
    $$
    We have $d(H,{\rm Id})<\epsilon$. 

    By the choice of $\tau_{\sigma}$, for all $x\in X\setminus L$,
    $$
    HTH^{-1}=S,
    $$
     thus, $d(HTH^{-1},S)\leq \mu(L)<\delta$.
\end{proof}
We continue the proof of $T_1\leftrightsquigarrow T_2$.  Fix an open neighborhood $\mathcal{V}$ of the identity in ${\rm Aut}(X,\mu)$ and $\mathcal{U}\subset {\rm Aut}(X,\mu)$ be any open set. We may assume that $\mathcal{V}$ contains an $\epsilon$-neighborhood of the identity.

Since every orbit is dense, we can find $H_1,H_2$ such that $H_iT_iH_i^{-1}\in \mathcal{U}$ for all $i\in \{1,2\}$ and 
$$
d(H_1T_1H_1^{-1},H_2T_2H_2^{-1})<\frac{\epsilon}{2}.
$$
Now we fix $\delta>0$. By Lemma \ref{keylemmaforeman}, applied to $T=H_1T_1H_1^{-1}$, $S=H_2T_2H_2^{-1}$ and $\mathcal{W}$  the  ${\rm min}\{\delta, \frac{\epsilon}{2}\}$-neighborhood of $H_2T_2H_2^{-1}$, we can find an $T'\in {\rm Aut}(X,\mu)$ such that $T'$ has the same orbits almost everywhere as $H_1T_1H_1^{-1}$ and 
$$
d(T',H_2T_2H_2^{-1})<{\rm min}\{\delta, \frac{\epsilon}{2}\}.$$ 
Thus, we have $d(T',H_1T_1H_1^{-1})<\epsilon$. By Lemma \ref{closeorbit} applied to $T=H_1T_1H_1^{-1}$ and $S=T'$, we can find an $H\in \mathcal{V}$ such that 
$$d(HH_1T_1H_1^{-1}H^{-1},T')< \delta.
$$ 
This implies that 
$$
d(HH_1T_1H_1^{-1}H^{-1},H_2T_2H_2^{-1})<2 \delta.
$$
As a result, we have 
$$
H_2T_2H_2^{-1}\in \overline{\mathcal{V}H_1T_1H_1^{-1}}.
$$
The proof of the other direction is symmetric. Take $g^{T_1}=H_1$ and $g^{T_2}=H_2$ in the definition of unbalancedness. We get $T_1 \leftrightsquigarrow T_2$.
\newline

\end{proof}

\section{Acknowledgments}
The author thanks François Le Ma\^{i}tre for valuable comments on a preliminary draft of the paper and Aristotelis Panagiotopoulos for helpful discussions on unbalancedness.
\bibliographystyle{plain} 
\bibliography{bibliography} 

\end{document}